# Principles and Guidelines for Evaluating Social Robot Navigation Algorithms

### v1

@@ VI-B2 Challenges Based on How Robots Change People
Beyond the novelty effect of the first encounters, people will refine their behavior around a robot as they interact with it for a longer period of time. A person will behave and react differently to a navigating robot on the tenth interaction than on the one-hundredth time [81, 82]. Answering questions regarding how people adapt to the presence of navigating robots remains mostly unanswered, and more research on this topic is needed [83], but studies of other social interactions such as asking favors [64] and information delivery [62] indicate that adaptation is likely.
@@ VI-B4 The Limitations of Social Metrics Themselves
    Human ratings are subjective. Human ratings are by their very nature subjective, and they depend on many factors such as cultural context, environmental context, goals or priorities within a scenario, or their overall familiarity with robots. It is important to account for the factors that all human participants experience, as well as attempt to characterize unique factors relevant to the scenario that can affect how they receive the scenario.
@@ VI-C Assessment of Existing Social Navigation Metrics
New metrics are often created to address issues that come up in new scenarios, and as social navigation is being deployed in increasingly many new environments, more metrics are created to address these scenarios. It is also unsurprising that new metrics will be of particular value in the environments that demanded their creation. This means that the number of metrics available to assess performance can be daunting, and their value very context-dependent.
@@
Number of collisions in the trajectory; when averaged it may be referred to as Collision Rate (CR). If one collision terminates the episode, CR is the complement of SR.
@@
Success weighted using the normalized inverse path length, i.e., weighted using path length divided by the maximum of the minimum distance and the path length [15].
@@ VI-E Metric Guidelines
    M6 - Set up experiments consistently to avoid bias: Environmental complexity, subject selection, robot familiarity, survey fatigue and differing experimental setups can all distort metrics. In particular, consider the use of well-designed scenarios (see Section VII) to make metrics easier to compare.
@@ VII Social Navigation Scenarios
  • Proponents of large-scale in-the-wild studies argue that these studies have good ecological validity, uncover long- tail behaviors, enable more reliable assessments of human perceptions of robot behavior, and enable data collection for unsupervised and reinforcement learning. These studies can have good statistical reliability and can generate large datasets; however, these studies are expensive to run, take a long time to generate answers, require heavyweight software architectures, and are suitable for policies that are already reliable.
  • Proponents of controlled in-the-lab scenarios argue these scenarios can also have good ecological validity, prevent regressions on known scenarios, enable scientific analysis of algorithms and behavior, and enable data collection for supervised and imitation learning. These studies are cheaper to run, generate data quickly, require less complex software, and are more appropriate for fast iteration on policies that are in an earlier stage of development; however, it is harder to uncover long-tail behaviors or to generate large datasets.
@@
In the following section, we outline a methodology for identifying and specifying social navigation scenarios which supports this breadth of usage, present a “Social Navigation Scenario Card” which enables scenarios to be clearly defined and disseminated, list common scenarios in the literature, present example scenario cards for some of the most commonly used ones, and conclude with guidelines for scenario development and usage.
@@ VII-A Scenario Design Methodology
To facilitate communicating scenarios, we propose a Social Navigation Scenario Card, presented next.
@@ VII-B Social Navigation Scenario Cards
To define a scenario, we propose a “Scenario Card” approach which labels the scenario with a set of features that unambiguously define it, having the following three major elements: (a) Scenario Metadata that define the name, description, and scientific purpose of the scenario; (b) Scenario Definition which clearly describes the environment, intended human behavior, and intended robot task; and (c) a Scenario Usage Guide, which provides additional information for specialized usages such as evaluation metrics, success and failure criteria.
@@ VII-B1 Scenario Metadata
The Scenario Metadata identifies a scenario in an unambiguous way for other researchers, including the type of the scenario (doorway, hallway, etc.), its name, its description, and its scientific purpose (crowd navigation, low-density pedestrian, interactive, etc.). For exampke, the common head-on pedestrian approach scenario might be labeled Frontal Approach, which we will use as a running example.
  • Scenario Type: Scenarios can be grouped into broad classes such as head-on approaches versus intersections, doorways and elevators, crowd versus group, interactive and accompanying, and so on. Identifying the group a scenario belongs to can help researchers decide whether to include a scenario for coverage or exclude it as redundant.
@@
  • Description: The scenario should have a brief description that communicates what is intended to happen in the scenario.
@@ VII-B2 Scenario Definition
  • Geometric Layout: Scenarios often occur in specific physical environments, such as corridors, doorways, at blind corners, or near elevators. The important features of the environment should be noted; features that can vary should also be noted.
@@ VII-B3 Scenario Usage Guide
  • Labeling Criteria: A clear set of criteria should be provided so that scenarios can be labeled in logs data or rejected in the event of a structured run. For example, for an intersection scenario, one could demand that the robot passes within two meters of the human and that their paths at least potentially cross.
@@ VII-C Example Social Navigation Scenarios
The columns of Table III capture many of the features of the Social Navigation Scenario card, though we cannot list all of them for space. Referring back to our running example, II shows an example of how the Social Navigation Scenario Card could be retroactively applied to one of the most common social navigation scenarios, Frontal Approach, which appears in [30, 6, 10] among others.
@@ VIII-D Properties of a Good Social Navigation Benchmark
    Guideline B1: Evaluate Social Behavior: A good benchmark should evaluate the properties of algorithms in social scenarios which involve humans and robots interacting. Therefore, a social benchmark should have metrics related to social behavior and not just contain pure navigation metrics such as Success weighted by Path Length [15] or pure task metrics such as success rates.
@@
(e) ucy
(f) eth
(g) epid
(h) sdd
(i) efl
(j) wildtrack
(k) crowdbot
(l) SocNav1
(m) SocNav2
(n) DynaBARN
(o) SACSoN
@@ IX Social Navigation Datasets
In this section, we provide a deeper look at datasets with regards to the factors listed in Sec. V. First we review desired dataset characteristics, noting that analyzing datasets require drilling in deeper on the existing factors, such as robot hardware, sensors, and behavior authoring methods, as well as additional factors for analysis such as data collected, dataset coverage, sampling distribution, annotations, and privacy and fairness handling. Then we use these factors to analyze several existing datasets, including jrdb [134], thor [135], Trajnet++ [136], ucy [125], eth [126], epid [137], sdd [138], efl [139], wildtrack [140], scand [141, 142], MuSoHu [143], CrowdBot [144], DynaBARN [23], SocNav1 [74], SocNav2 [75], and SACSon [145], reviewing them with respect to the criteria (Fig. 9).
@@ IX-A Expanding the Factors for Dataset Analysis
As social navigation datasets contain humans, privacy is a big concern. Decisions must be made whether to anonymize humans and how to comply with privacy protection regulations.
@@ IX-B Existing Social Navigation Datasets
Benchmark Dataset THOR SCAND UCY/ETH Trajnet++ EPID SocNav1 SocNav2
@@
Benchmark Dataset SDD EFL LCAS WILDTRACK JRDB CrowdBot DynaBARN SACSoN
@@
Autonomous policy interacting with humans.
@@
scand [142, 141] is a public dataset providing socially compliant navigation demonstrations recorded via teleoperating two different mobile robots in a socially compliant manner by human demonstrators. The objective behind the scand dataset is to study the social navigation behavior of robots in the presence of human crowds. Similar to scand, MuSohu [143] includes 3D lidar scans, RGBD camera images, 360°  camera images, IMU data, and ambient sound collected from a sensor suite mounted on a helmet worn by humans walking around public spaces (instead of on a teleoperated robot), from which social robot navigation can be learned. The portable sensor suite allows social human navigation data to be collected in the wild with a low setup cost, making MuSoHu easily extendable.
Also similar to scand, lcas [147] is a public dataset containing 3D lidar scans collected using a mobile robot teleoperated in heavily crowded environments. However, unlike scand, the robot is not necessarily teleoperated in a socially compliant manner. The focus of the lcas dataset is to solve perception-related challenges in social navigation, such as online human detection.
The ucy/eth [125, 126] dataset consists of human trajectories recorded in public spaces from a bird’s eye view vantage point using an RGB camera. The trajectories are extracted by tracking humans from the bird’s eye view images. The motivation behind the ucy dataset is to provide real-world trajectories of humans navigating among other humans in the scene so one can replicate-by-copying such trajectories in a simulator. Trajectories from the ucy dataset can be used to simulate a diverse set of realistic social scenarios. Pellegrini et al.[4] propose conditioning the predicted future trajectory also on scene knowledge and social interactions among agents.
@@
epid [137] is again similar to eth and ucy, while providing a much higher number of humans captured in the dataset, the camera is fixed overhead roughly about 23 meters from the floor. Humans are detected by processing this bird’s eye view image from the scene and tracking them in the scene.
Stanford Drone Dataset (sdd) [138] is similar to ucy and eth since it also provides a bird’s eye view frame, recorded using a drone (unlike ucy and eth that use a statically mounted camera). Compared to ucy and eth, the unique selling point of this dataset is large-scale images and videos of diverse scenarios including bicyclists, skateboarders, cars, buses, and golf carts navigating in the real world.
Egocentric Future Localization (efl) [139] provides RGBD sequences of frames from the perspective of a human, walking in different scenes. This dataset includes various indoor and outdoor scenes such as Parks, Malls, and a Campus, with various activities such as walking, shopping, and social interaction. The main focus of this work seems to be on human trajectory prediction in novel scenes. Note that the website hosting this dataset is no longer accessible.
wildtrack [140] is similar to ucy and eth. A GoPro camera is mounted in an outdoor environment scene consisting of crowds of people walking around. This dataset focuses on person detection in the presence of severe obstacles such as other humans and static obstacles in the scene.
@@ X Simulation-based Evaluation
The fundamental requirement for a social navigation simulator is the ability to simulate two agents at one time in a social encounter - without that, it’s just traditional navigation. Beyond this core requirement, social navigation simulators span the gamut from supporting crowds of simplified agents to test dynamic navigation algorithms to simulators that recreate human appearances, footsteps, behavioral diversity, and environmental interactivity. Most benchmarks discussed in Section VIII rely on a simulator to make benchmarking efficient, repeatable and scalable.
@@ X-A Expanding the Factors for Simulator Analysis
Pedestrians can move on pre-recorded trajectories without reacting to other agents, or may react using a model such as the Social Forces Model (SFM) [111] or ORCA [122]. Pedestrian behavior may be also modulated with individual attitudes, behavioral styles or social activities specified by higher-level modules.
@@ X-B Existing Social Navigation Simulators
CrowdNav [152] is a 2D simulator for multi-agent scenarios using ORCA [153, 127] to orchestrate pedestrian discs around policy-controlled discs in simplified environments.
@@
The HuNavSim [25] benchmark contains a simulator using Behavior Trees (BT) [124] and the Social Forces Model (SFM) [111] to provide a variety of human behaviors ranging from indifferent, surprised, curious, fearful and aggressive. HuNavSim can work with various simulators, and represents both human gait and robot morphologies.
@@ X-C3 Pedestrian Reactivity
Motion models for pedestrians such as the Social Forces Model (SFM) [111] and Optimal Reciprocal Collision Avoidance (ORCA) [122] enable them to move in reaction to changes in the environment. While no model of human motion is perfect, the modeling of reactive agents in simulation allows researchers to explore how changes made to the environment by different robot policies affect task performance.

### v4

@@ VI-B2 Challenges Based on How Robots Change People
Beyond the novelty effect of first encounters, people will refine their behavior around a robot as they interact with it over time. A person will behave and react differently to a navigating robot on the tenth interaction than the hundredth [81, 82]. More research is needed on how people adapt to the presence of navigating robots [83], but studies of other social interactions such as asking favors [64] and information delivery [62] indicate that adaptation is likely.
@@ VI-B4 The Limitations of Social Metrics Themselves
    Human ratings are subjective. Human ratings are by their very nature subjective, and they depend on many factors such as cultural context, environmental context, goals or priorities within a scenario, or their overall familiarity with robots. It is important to account for the factors that all human participants experience, as well as attempt to characterize unique factors relevant to the scenario that can affect how they perceive the scenario.
@@ VI-C Assessment of Existing Social Navigation Metrics
New metrics are often created to address issues that come up in new scenarios, and as social navigation is being deployed in increasingly many new environments, more metrics are being created to address these scenarios. It is also unsurprising that new metrics will be of particular value in the environments that demanded their creation. This means that the number of metrics available to assess performance can be daunting, and their value very context-dependent.
@@
Number of collisions in the trajectory. When averaged it is referred to as Collision Rate (CR).
@@
Success weighted using normalized inverse path length, i.e., weighted using path length divided by the max of the min distance and path length [15].
@@ VI-E Metric Guidelines
    M6 - Set up experiments consistently to avoid bias: Environmental complexity, subject selection, robot familiarity, survey fatigue and differing experimental setups can all distort metrics. Use well-designed scenarios (see Section VII) to make metrics easier to compare.
@@ VII Social Navigation Scenarios
  • Proponents of large-scale in-the-wild studies argue they have good ecological validity, uncover long-tail behaviors, enable more reliable assessments of human perceptions of robot behavior, and enable data collection for unsupervised and reinforcement learning. These studies can have good statistical reliability and can generate large datasets; however, they are expensive, time-consuming, require heavyweight software architectures, and are suitable for policies that are already reliable.
  • Proponents of controlled in-the-lab scenarios argue they can also have good ecological validity, prevent regressions on known issues, enable scientific analysis of algorithms and behavior, and enable data collection for supervised and imitation learning. These studies are cheaper to run, generate data quickly, require less complex software, and are more appropriate for iterating on policies in an earlier stage of development; however, it is harder to uncover long-tail behaviors or to generate large datasets.
@@
In the next section, we outline a methodology for identifying and specifying social navigation scenarios which supports this breadth of usage, present a “Social Navigation Scenario Card” which enables scenarios to be clearly defined and disseminated, list common scenarios in the literature, and conclude with guidelines for scenario development and usage.
@@ VII-A Scenario Design Methodology
Social navigation scenario development can be seen as a step towards the more formal scenarios engineering approach being adopted in intelligent vehicles research [106, 107, 108]. To facilitate communicating scenarios, we propose a Social Navigation Scenario Card, presented next.
@@ VII-B Social Navigation Scenario Cards
Following work on “model cards” in the machine learning community [110], we propose a “Scenario Card” approach to defining scenarios which which labels the scenario with a set of features that unambiguously define it. Scenario cards have the following three major elements: (a) Scenario Metadata that define the name, description, and scientific purpose of the scenario; (b) Scenario Definition which clearly describes the environment, intended human behavior, and intended robot task; and (c) a Scenario Usage Guide, which provides additional information for specialized usages such as evaluation metrics, success and failure criteria.
@@ VII-B1 Scenario Metadata
The Scenario Metadata identifies a scenario in an unambiguous way for other researchers, including the type of the scenario (doorway, hallway, etc.), its name, its description, and its scientific purpose (crowd navigation, low-density pedestrian, interactive, etc.). For example, a head-on pedestrian approach scenario might be labeled Frontal Approach, which we will use as a running example.
  • Scenario Type: Scenarios can be grouped into broad classes such as head-on approaches versus intersections, doorways and elevators, crowd versus group, interactive and accompanying, and so on. Identifying the group a scenario belongs to can help researchers decide whether to include it for coverage or exclude it as redundant.
@@
  • Description: The scenario should have a brief description that communicates what is intended to happen in it.
@@ VII-B2 Scenario Definition
  • Geometric Layout: Scenarios often occur in specific physical environments, such as corridors, doorways, at blind corners, or near elevators. The important features of the environment should be noted; features that can vary should also be noted so the scenario is not overspecified.
@@ VII-B3 Scenario Usage Guide
  • Labeling Criteria: A clear set of criteria should be provided so that scenarios can be labeled in logs data or rejected in the event of a structured run. For example, for an intersection scenario, one could demand that the robot passes within two meters of the human and that their intended paths at least potentially cross.
@@ VII-C Example Social Navigation Scenarios
The columns of Table III capture many of the features of the Social Navigation Scenario card, though we cannot list all of them for space. Referring back to our running example, Table II shows an example of how the Social Navigation Scenario Card could be retroactively applied to one of the most common social navigation scenarios, Frontal Approach, which appears in [30, 6, 10] among others.
@@ VIII-D Properties of a Good Social Navigation Benchmark
    Guideline B1: Evaluate Social Behavior: A good social benchmark should evaluate the properties of algorithms in social scenarios which involve humans and robots interacting. Therefore, a social benchmark should have metrics related to social behavior and not just contain pure navigation metrics such as Success weighted by Path Length [15] or pure task metrics such as success rates.
@@
(e) eifpd
(f) sdd
(g) efl
(h) wildtrack
(i) crowdbot
(j) SocNav1
(k) SocNav2
(l) DynaBARN
(m) SACSoN
(n) SG-LSTM
@@ IX Social Navigation Datasets
In this section, we provide a deeper look at datasets with regards to the factors listed in Sec. V. First we review desired dataset characteristics, noting that analyzing datasets requires drilling in deeper on factors such as robot hardware, sensors, and behavior authoring methods, as well as additional factors for analysis such as data collected, dataset coverage, sampling distribution, annotations, and privacy and fairness handling. Then we use these factors to analyze several datasets, including jrdb [138], thor [139], Trajnet++ [140], eth/ucy [129, 130], eifpd [141], sdd [142], efl [143], wildtrack [144], scand [145, 146], MuSoHu [147], CrowdBot [148], DynaBARN [23], SocNav1 [74], SocNav2 [75], SACSon [149], and sg-lstm [150], reviewing them with respect to the criteria (Fig. 9).
@@ IX-A Expanding the Factors for Dataset Analysis
As social navigation datasets contain humans, privacy is an important concern. Decisions must be made whether to anonymize humans and how to comply with privacy protection regulations.
@@ IX-B Existing Social Navigation Datasets
Benchmark Dataset THOR SCAND ETH/UCY Trajnet++ EIFPD SocNav1 SocNav2
@@
Benchmark Dataset SDD EFL LCAS WILDTRACK JRDB Crowd-Bot Dyna-BARN SACSoN SG-LSTM
@@
Autonomous policy interacting with humans
Curated interaction, movements and formation of pedestrian groups
@@
Color and depth frames, pedestrian and group bounding boxes
@@
Paths, green spaces, study spaces, cafes, gatherings, weather events
@@
Group size, walking speeds, proximity, cohesiveness, interactions.
@@
scand [146, 145] is a public dataset providing socially compliant navigation demonstrations recorded via teleoperating two different mobile robots in a socially compliant manner by human demonstrators. The objective behind the scand dataset is to study the social navigation behavior of robots in the presence of human crowds. Similar to scand, MuSohu [147] includes 3D lidar scans, RGBD camera images, 360°  camera images, IMU data, and ambient sound collected from a sensor suite mounted on a helmet worn by humans walking around public spaces (instead of on a teleoperated robot), from which social robot navigation can be learned. This allows social human navigation data to be collected in the wild with a low setup cost, making MuSoHu easily extendable.
Also like scand, lcas [152] is a public dataset containing 3D lidar scans collected using a mobile robot teleoperated in heavily crowded environments. Unlike scand, the robot is not necessarily teleoperated in a socially compliant manner. The focus of lcas is to solve perception-related challenges in social navigation, such as online human detection.
The eth/ucy [129, 130] dataset consists of human trajectories recorded in public spaces from a bird’s eye view vantage point using an RGB camera. The trajectories are extracted by tracking humans from the bird’s eye view images. The motivation behind the eth/ucy dataset is to provide real-world trajectories of humans navigating among other humans in the scene so one can replicate-by-copying such trajectories in a simulator. Trajectories from the eth/ucy dataset can be used to simulate a diverse set of realistic social scenarios. [130] propose conditioning the predicted future trajectory also on scene knowledge and social interactions among agents.
@@
The Ediburgh Informatics Forum Pedestrian Database (eifpd) [141] is again similar to eth/ucy, while providing a much higher number of humans captured in the dataset, the camera is fixed overhead roughly about 23 meters from the floor. Humans are detected by processing this bird’s eye view image from the scene and tracking them in the scene.
Stanford Drone Dataset (sdd) [142] is similar to eth/ucy since it also provides a bird’s eye view frame, recorded using a drone (unlike eth/ucy that uses a statically mounted camera). Compared to eth/ucy, the unique selling point of this dataset is large-scale images and videos of diverse scenarios including bicyclists, skateboarders, cars, buses, and golf carts navigating in the real world.
Egocentric Future Localization (efl) [143] provides RGBD sequences of frames from the perspective of a human in various indoor and outdoor scenes such as Parks, Malls, and a Campus, with various activities such as walking, shopping, and social interaction. efl’s focus is human trajectory prediction in novel scenes.
wildtrack [144] is similar to eth/ucy. A GoPro camera is mounted in an outdoor environment scene consisting of crowds of people walking around. This dataset focuses on person detection in the presence of severe obstacles such as other humans and static obstacles in the scene.
@@
The sg-lstm [150] dataset is a curated dataset designed to provide insight into pedestrian behavior collected on Purdue University’s West Lafayette, Indiana campus using a GO1 Edu robot by Unitree Robotics navigating unobtrusively among pedestrians. SG-LSTM focuses on the interactions, movements, and group formations of pedestrians in a variety of scenarios including campus thoroughfares, gatherings, dining and study areas, green spaces, and inclement weather events.
@@ X Simulation-based Evaluation
The fundamental requirement for a social navigation simulator is the ability to simulate two agents at one time in a social encounter - without that, it’s just traditional navigation. Beyond this core requirement, social navigation simulators span the gamut from supporting crowds of simplified agents that test dynamic navigation algorithms to simulators that recreate human appearances, footsteps, behavioral diversity, and environmental interactivity. Most benchmarks discussed in Section VIII rely on a simulator to make benchmarking efficient, repeatable and scalable.
@@ X-A Expanding the Factors for Simulator Analysis
Pedestrians can move on pre-recorded trajectories without reacting to other agents, or may react using a model such as the Social Forces Model (SFM) [115] or ORCA [126]. Pedestrian behavior may be also modulated with individual attitudes, behavioral styles or social activities specified by higher-level modules or using techniques such as behavior trees [128].
@@ X-B Existing Social Navigation Simulators
CrowdNav [157] is a 2D simulator for multi-agent scenarios using ORCA to orchestrate pedestrian discs around policy-controlled discs in simplified environments.
@@
The HuNavSim [25] benchmark contains a simulator using SFM and behavior trees to provide a variety of human behaviors ranging from indifferent, surprised, curious, fearful and aggressive. It can work with various simulators, and represents both human gait and robot morphologies.
@@ X-C3 Pedestrian Reactivity
Motion models for pedestrians such as SFM [115] and ORCA [126] enable them to move in reaction to changes in the environment. While no model of human motion is perfect, the modeling of reactive agents in simulation allows researchers to explore how changes made to the environment by different robot policies affect task performance.